the test statistic of z = 1.39 is obtained when testing the claim that p>0.8.  

a. Identify the hypothesis test as being two tailed, left tailed, or right tailed.

b. Find the P-value.

c. Using a significance level alpha = 0.05, should we reject Ho or should we fail to reject Ho?

d. What is the correct conclusion for this test?
Answer #1

Solution :

Given that

a ) This is the riight tailed test .

The null and alternative hypothesis is

H0 : p = 0.8

Ha : p >0.8

Test statistic = z = 1.39

P(z > 1.39 ) = 1 - P(z < 1.39 ) = 1 - 0.9177

b ) P-value = 0.0823

c ) = 0.05

P-value >

0.0823 > 0.05

d ) Fail to reject the null hypothesis .

There is insufficient evidence to suggest that